You lend $1,128 to a friend and he agrees to pay you back next year plus he will pay you interest of 7.93%. If the inflation rate is 2.41%, then in real terms how much extra money will you have after your friend pays you back? Enter a number rounded to two decimal places with no other characters.
Answer #1

Current real value of money = 1128

Total money you get after lending = Principal + interest

= 1128 + 0.0793*1128

= 1217.45

Real income = income/ (1+inflation) = 1217.45/1.0241 = 1188.79

Thus, additional money after your friend pays back

= 1188.79 - 1128 = 60.79